# SwapAnything: Enabling Arbitrary Object Swapping in Personalized Visual Editing

### v1

@@ Abstract
Figure 1: SwapAnything results on various personalized image swapping tasks. SwapAnything is adept at precise, arbitrary object replacement in a source image with a personalized reference, and achieves high-fidelity swapping results without influencing any context pixels. We demonstrate its general effect in single-object, multi-object, partial-object, and cross-domain swapping tasks.
@@ 1 Introduction
In today’s digital age marked by the prolific creation and widespread sharing of personal content, generative models (Rombach et al., 2022; Crowson et al., 2022; Ding et al., 2021; Blattmann et al., 2022) have risen as a potent tool for self-expression, storytelling, and amplifying the impact of visual narratives. Among existing image editing methods (Meng et al., 2022; Hertz et al., 2022; Chen et al., 2023b) for content creation that empowered individuals to convey their creative instincts, weave captivating narratives into their visual stories, and enhance the quality of their visual representations, personalize content swapping (Gu et al., 2023a; Li et al., 2023b; Gu et al., 2023b), which targets at creating and compositing new images with user-specified visual concept, attracted significant interest due to its wide-ranging applications in E-commerce, entertainment, and professional editing.
@@
Existing works often fall short of addressing these challenges. Most of existing research (Ruiz et al., 2023; Shi et al., 2023; Chen et al., 2023b; Jia et al., 2023; Wang et al., 2024) are focused on personalized image synthesis, which seeks to create new images with personalized content. Although these approaches can synthesize high-fidelity content, they cannot edit or replace the visual content in an existing image.  Li et al. (2023b); Cao et al. (2023); Meng et al. (2022); Avrahami et al. (2023) have tried to remove and regenerate the object via masks, they often struggle to adapt the new concept into the image. Recently studies (Patashnik et al., 2023; Gu et al., 2023a; Tumanyan et al., 2023) focus on object swapping and replacement by tweaking intermediate variables affecting the object’s features. However, this approach lacks the precision needed for localized object swapping, resulting in limited visual qualities. Besides, those methods mainly work on main subject swapping, and they are incapable of arbitrary object swapping.
@@ 2.1 Text-guided Image Editing
With recent progress in diffusion model, text-guided Image editing has been largely explored (Avrahami et al., 2023). Image editing driven by text aims to alter an existing image following the textual guidelines provided, while ensuring certain elements or features of the original image remain unchanged. Initial efforts using GAN models (Karras et al., 2019) were restricted to specific object domains. However, diffusion-oriented techniques (Zhang et al., 2023; Nichol et al., 2022; Feng et al., 2023) have surpassed this limitation, enabling text-driven image modifications. While these techniques produce captivating outcomes, many grapple with localized edits. As a result, manual masks (Meng et al., 2022; Zeng et al., 2023; Meng et al., 2022) are often help to delineate the editing zones. Though employing cross-attention (Hertz et al., 2022) or spatial attributes (Tumanyan et al., 2023) can achieve localized edits, they face challenges with flexible changes (like altering poses) and preserving the initial image composition.
@@ 2.2 Object Driven Image Editing
In object-driven image generation, the object could be inverted into the texual space so to represented by a new token such as “***”  (Gal et al., 2023; Ruiz et al., 2023; Chen et al., 2023a; Shi et al., 2023; Tewel et al., 2023). Image editing guided by exemplars spans a vast array of uses. Much of the research in this area (Wang et al., 2019; Huang et al., 2018; Zhou et al., 2021) falls under the umbrella of example-based image transformation tasks. These tasks leverage various sources of information, from stylized images (Liu et al., 2021; Deng et al., 2022; Zhang et al., 2022) and layouts (Yang et al., 2023b; Li et al., 2023c; Jahn et al., 2021), to skeletons (Li et al., 2023c) and sketches or outlines (Seo et al., 2023). Given the versatility of stylized images, the art of image style adaptation (Liao et al., 2017; Zhang et al., 2020) has garnered significant interest. These methods primarily focus on establishing a detailed match between the input and reference visuals but falter when it comes to localized alterations. To facilitate localized changes, especially with flexible transformations, tools like bounding boxes and skeletons have been introduced. However, these tools often demand manual input, which can be challenging for users. A novel approach (Yang et al., 2023a) frames exemplar-driven image editing as a task of filling in gaps, maintaining the context while transferring semantic elements from the reference to the original image. While DreamEdit (Li et al., 2023b) employs a step-by-step gap-filling method for object substitution, it doesn’t effectively bridge the connection between the original and desired objects. In contrast, our method ensures that crucial features, like body movements and facial expressions, stay consistent. Meanwhile, Photoswap (Gu et al., 2023a), CustomEdit (Choi et al., 2023) and BLIP-Diffusion (Li et al., 2023a) achieved personalized object swapping. However, it does not keep the non-object pixel background intact. While our method directly focuses on the editing area without influencing other objects and background.
@@ 3 Preliminary
In our research, we employ the pre-trained text-to-image diffusion framework Stable Diffusion Rombach et al. (2022). This model encodes images into a latent space and incrementally denoises the encoded latent representation. The Stable Diffusion model operates on a U-Net architecture (Ronneberger et al., 2015), where the latent representation ${\boldsymbol{z}}_{t-1}$ at any given step is derived from the text prompt $P$ and the previous latent state ${\boldsymbol{z}}_{t}$, as indicated by the following equation:
@@
Figure 2: Overview of SwapAnything on swapping a object from a source image ($I_{src}$) into a personalized concept (<*><{*}>< * >) to get the target image ($I_{target}$). The personalized concept is first converted into textual space to be treated as concept appearance. Meanwhile, the source image is first inverted into initial noise to obtain U-Net variables (including latent feature, attention map, and attention output). Targeted variable swapping preserves the context pixels in the source image. The appearance adaptation process then utilizes these informative variables to integrate the concept into the target image.
@@ 4.1 Targeted Variable Swapping
Intermediate variables in the U-Net of a diffusion model have been proven informative about the content of the generated image (Gu et al., 2023a; Hertz et al., 2022; Cao et al., 2023; Tumanyan et al., 2023). They usually focus on the study of variables inside of U-net structure such as attention map, and attention output, while the output of U-Net at each diffusion step, i.e., latent image feature ${\boldsymbol{z}}$ is not widely explored before. We argue that the latent image feature ${\boldsymbol{z}}$ contains more information on image content control. The image generation process for the latent diffusion model is achieved by denoising the ${\boldsymbol{z}}$ to arrive at a clear representation of a high-quality image, whereas all other variables inside of U-net indirectly affect the image by impacting ${\boldsymbol{z}}$. In contrast to simply swapping ${\boldsymbol{z}}$ like other variables, which would erase the new image’s unique details and result in a mere duplication of the original image, our investigation revealed a significant correlation between the latent feature ${\boldsymbol{z}}$ and the produced image, including a pixel-level correspondence. In the left part of Fig. 3, we visualize the main component of the averaged latent feature $z$ across all diffusion steps. The finding that It has a part-to-part correspondence with the generated images indicates the potential of localized editing by manipulating the latent feature.
Consequently, we consider a strategy where only the context pixels within ${\boldsymbol{z}}$ are altered, affecting solely the intended pixel. Although this method yields suboptimal outcomes, akin to merging two unrelated images, it prompts us to devise two remedial measures. Firstly, we suggest limiting the exchange of the latent feature to the initial stages of diffusion, allowing subsequent steps to smooth out any discordance in the latent space. Furthermore, our exploration into Unet’s cross-attention map ${\boldsymbol{M}}$, self-attention map ${\boldsymbol{A}}$, and self-attention output ${\boldsymbol{\psi}}$ reveal their potential in mitigating artifacts. Swapping those facilitates the alignment of the latent features between the source and target images before the partial-swapping between them. In short, all the variables mentioned above in both the source image and target image generation process could be resized into the shape of the mask, where the mask can be utilized for the swapping process.
@@
Here $V$ includes latent feature $z$, and other assistant variable cross-attention map ${\boldsymbol{M}}$, self-attention map ${\boldsymbol{A}}$, and self-attention output ${\boldsymbol{\phi}}$. $f(\cdot)$ means the transformation process to the shape of the mask, while $g(\cdot)$ means the transformation back to the original space. For simplicity, we ignore all $f(\cdot)$ and $g(\cdot)$ in the following text. The content in the latent feature of the source image is changing as the diffusion process continues. Therefore, the location of the correspondent pixel in latent space may change over diffusion steps. A direct solution is to decode the latent feature ${\boldsymbol{z}}$ into an image at each step and extract the mask dynamically according to the object location in the generated image. However, we find that a changing mask usually confuses the model and leads to a less optimal performance. Therefore, while using the same high-quality mask through the diffusion process. We find that the mask could either be extracted from the source image directly using an off-the-shelf model or from the generation process as in Simsar et al. (2023); Patashnik et al. (2023). Please check the Appendix for more details.
@@ 4.2.2 Style Adaptation
An ideal object swapping should keep the style unchanged. The object information in the generated variables is injected via the new concept token. Some style attributes could be already bound with the token. Therefore, solely generating the foreground information via the text prompt might lead to style inconsistency. Recently, Karras et al. (2019); Park et al. (2019) found that adding such normalization layers can help improve the conditional image generation quality because such activation functions modulation. Unlike from them, in our work, we employ the AdaIN (Adaptive Instance Normalization) to modulate the swapping features with spatial constraints. We follow Eq. 6 and Eq. 7 to denormalize the $V_{concept}$ with the mean and variance from $V_{src}$ in each time step for $V_{target}$ during the image generation process. As a result, we find that by modulating the concept feature, the generated content can adaptively follow the original style in the source image.
@@ 4.2.3 Scale Adaptation
Guidance from an external classifier in the inference process of Diffusion models could influence the diffusion noise to control the generated image. Epstein et al. (2023) shows that the guidance can also be used on the attention map to control the generation. Similarly, we adapt the mask guidance Eq. 9 to better align the shape between the source object and the target object.
@@
where $s$ is the classifier-free guidance strength and $v$ is an additional guidance weight for $g$. As with classifier guidance, we scale by $\sigma_{t}$ to convert the score function to a prediction of $\mathbf{\epsilon}_{t}$. $Shape(\mathbf{M_{src}}{})$ denotes the object shape as identified in the attention layer. Here the energy function $g$ is set as $\|\mathbf{M_{src}}{}-Shape(\mathbf{M_{src}}{})\|_{1}$ to calculate the shape difference between the original object mask and the extracted shape in the attention layer, which indicates the deviation between the ideal shape and shape during the diffusion process.
@@ 4.2.4 Content Adaptation
the dilation of the mask $\mathbf{M_{src}}{}$ using the structuring element K, where denotes the dilation operation and G is the Gaussian kernel. The asterisk *** denotes the convolution operation. $S^{\prime}$ is the final soft mask.
@@ 5.1 Implementation Details
Here we introduce implementation details. In our paper, we used Stable Diffusion 2.1 as the pre-trained text-to-image diffusion model. DreamBooth (Ruiz et al., 2023) is used to convert the concept into textual space. We used null-text inversion (Mokady et al., 2023) based on DDIM inversion (Song et al., 2020) to boosts the accuracy and reliability of the inversion.
For object mask, we first detect the object with Grounding DINO (Liu et al., 2023) and then extract the mask using Segment Anything (Kirillov et al., 2023). For the targeting variable swapping process, we do 30 for latent image feature $z$, 20 steps for cross-attention map, 25 for the self-attention map, and 10 for the self-attention output, We conduct swapping in all U-Net layer. There is no additional operation for single-object, partial object, cross domain swapping. Multi-object swapping is achieved by conducting swapping operation on the previous swapped image.
Dataset. We conducted experiments on both human and non-human objects. For human swapping, we collect celebrities from internet searches. We employed the search prompt: “a photo of $<target>$”, where $<target>$ is the celebrity name. We collected images of 15 celebrities for the concept learning process. We also collected 500 images containing 1 or more people as the source images. For non-human object, we include DreamEdit (Li et al., 2023b) dataset and more concepts and its corresponding source images from Google search. In total, we aggregated 1,000 images. We will also release the collected dataset.
Baselines setting. Photoswap(Gu et al., 2023a), P2P(Hertz et al., 2022), PnP(Tumanyan et al., 2023), MasaCtrl(Cao et al., 2023) are attention variable based image editing methods, which are also compatible with our proposed Masked Latent Blending in Sec. 4.1 and Location Adaptation in Sec. 4.2. Therefore we boost their performance with those additional components. Without those, their performance would be much worse. Please see appendix for comparisons with their original methods. Besides, we use external mask to help the inpainting process in DreamEdit. SwapAnything is also compared with BlipDiffusion (Li et al., 2023a). Photoswap, P2P, PnP, and MasaCtrl, DreamEdit were equipped with the same DreamBooth model to grasp the new concept. Note that this would also indirectly include comparison with CustomEdit (Choi et al., 2023), since it also achieved personalized object swapping via equiping P2P with concept learning. CopyPaste involves directly transplanting the concept object in the concept image into the source object’s position.

### v3

# https://swap-anything.github.io/

@@ Abstract
Figure 1: SwapAnything results on various personalized image swapping tasks. SwapAnything is adept at precise, arbitrary object replacement in a source image with a personalized reference, and achieves high-fidelity swapping results without influencing any context pixels. We demonstrate its general swapping effect in single-object, multi-object, partial-object, and cross-domain swapping tasks.
@@ 1 Introduction
In today’s digital age marked by the prolific creation and widespread sharing of personal content, generative models [32, 7, 9, 2] have risen as a potent tool for self-expression, storytelling, and amplifying the impact of visual narratives. Among existing image editing methods [27, 15, 5] for content creation that empowered individuals to convey their creative instincts, weave captivating narratives into their visual stories, and enhance the quality of their visual representations, personalize content swapping [13, 22, 14], which targets at creating and compositing new images with user-specified visual concept, attracted significant interest due to its wide-ranging applications in E-commerce, entertainment, and professional editing.
@@
Existing works often fall short of addressing these challenges. Most of existing research [34, 36, 5, 18, 42] are focused on personalized image synthesis, which seeks to create new images with personalized content. Although these approaches can synthesize high-fidelity content, they cannot edit or replace the visual content in an existing image.  [22, 3, 27, 1] have tried to remove and regenerate the object via masks, they often struggle to adapt the new concept into the image. Recently studies [31, 13, 40] focus on object swapping and replacement by tweaking intermediate variables affecting the object’s features. However, this approach lacks the precision needed for localized object swapping, resulting in limited visual qualities. Besides, those methods mainly work on main subject swapping, and they are incapable of arbitrary object swapping.
@@ 2.1 Text-guided Image Editing
With recent progress in diffusion model, text-guided Image editing has been largely explored [1]. Image editing driven by text aims to alter an existing image following the textual guidelines provided, while ensuring certain elements or features of the original image remain unchanged. Initial efforts using GAN models [19] were restricted to specific object domains. However, diffusion-oriented techniques [46, 29, 11] have surpassed this limitation, enabling text-driven image modifications. While these techniques produce captivating outcomes, many grapple with localized edits. As a result, manual masks [27, 45, 27] are often help to delineate the editing zones. Though employing cross-attention [15] or spatial attributes [40] can achieve localized edits, they face challenges with flexible changes (like altering poses) and preserving the initial image composition.
@@ 2.2 Object Driven Image Editing
In object-driven image generation, the object could be inverted into the texual space so to represented by a new token such as “*”  [12, 34, 4, 36, 39]. Image editing guided by exemplars spans a vast array of uses. Much of the research in this area [41, 16, 49] falls under the umbrella of example-based image transformation tasks. These tasks leverage various sources of information, from stylized images [26, 8, 48] and layouts [44, 23, 17], to skeletons [23] and sketches or outlines [35]. Given the versatility of stylized images, the art of image style adaptation [24, 47] has garnered significant interest. These methods primarily focus on establishing a detailed match between the input and reference visuals but falter when it comes to localized alterations. To facilitate localized changes, especially with flexible transformations, tools like bounding boxes and skeletons have been introduced. However, these tools often demand manual input, which can be challenging for users. A novel approach [43] frames exemplar-driven image editing as a task of filling in gaps, maintaining the context while transferring semantic elements from the reference to the original image. While DreamEdit [22] employs a step-by-step gap-filling method for object substitution, it doesn’t effectively bridge the connection between the original and desired objects. In contrast, our method ensures that crucial features, like body movements and facial expressions, stay consistent. Meanwhile, Photoswap [13], CustomEdit [6] and BLIP-Diffusion [21] achieved personalized object swapping. However, it does not keep the non-object pixel background intact. While our method directly focuses on the editing area without influencing other objects and background.
@@ 3 Preliminary
In our research, we employ the pre-trained text-to-image diffusion framework Stable Diffusion [32]. This model encodes images into a latent space and incrementally denoises the encoded latent representation. The Stable Diffusion model operates on a U-Net architecture [33], where the latent representation ${\boldsymbol{z}}_{t-1}$ at any given step is derived from the text prompt $P$ and the previous latent state ${\boldsymbol{z}}_{t}$, as indicated by the following equation:
@@
Figure 2: Overview of SwapAnything on swapping a object from a source image ($I_{src}$) into a personalized concept (<><{*}>< ∗ >) to get the target image ($I_{target}$). The personalized concept is first converted into textual space to be treated as concept appearance. Meanwhile, the source image is first inverted into initial noise to obtain U-Net variables (including latent feature, attention map, and attention output). Targeted variable swapping preserves the context pixels in the source image. The appearance adaptation process then utilizes these informative variables to integrate the concept into the target image.
@@ 4.1 Targeted Variable Swapping
Intermediate variables in the U-Net of a diffusion model have been proven informative about the content of the generated image [13, 15, 3, 40]. They usually focus on the study of variables inside of U-net structure such as attention map, and attention output, while the output of U-Net at each diffusion step, i.e., latent image feature ${\boldsymbol{z}}$ is not widely explored before. We argue that the latent image feature ${\boldsymbol{z}}$ contains more information on image content control. The image generation process for the latent diffusion model is achieved by denoising the ${\boldsymbol{z}}$ to arrive at a clear representation of a high-quality image, whereas all other variables inside of U-net indirectly affect the image by impacting ${\boldsymbol{z}}$. In contrast to simply swapping ${\boldsymbol{z}}$ like other variables, which would erase the new image’s unique details and result in a mere duplication of the original image, our investigation revealed a significant correlation between the latent feature ${\boldsymbol{z}}$ and the produced image, including a pixel-level correspondence. In the left part of Fig. 3, we visualize the main component of the averaged latent feature $z$ across all diffusion steps. The finding that It has a part-to-part correspondence with the generated images indicates the potential of localized editing by manipulating the latent feature.
Consequently, we consider a strategy where only the context pixels within ${\boldsymbol{z}}$ are altered, affecting solely the intended pixel. Although this method yields suboptimal outcomes, akin to merging two unrelated images, it prompts us to devise two remedial measures. Firstly, we suggest limiting the exchange of the latent feature to the initial stages of diffusion, allowing subsequent steps to smooth out any discordance in the latent space. Furthermore, our exploration into Unet’s cross-attention map ${\boldsymbol{M}}$, self-attention map ${\boldsymbol{A}}$, and self-attention output ${\boldsymbol{\psi}}$ reveal their potential in mitigating artifacts. Swapping those facilitates the alignment of the latent features between the source and target images before the partial-swapping between them. In short, all the variables mentioned above in both the source image and target image generation process could be resized into the shape of the mask, where the mask can be utilized for the swapping process to get target variable:
@@
Here $V$ includes latent feature $z$, and other assistant variable cross-attention map ${\boldsymbol{M}}$, self-attention map ${\boldsymbol{A}}$, and self-attention output ${\boldsymbol{\phi}}$. $f(\cdot)$ means the transformation process to the shape of the mask, while $g(\cdot)$ means the transformation back to the original space. For simplicity, we ignore all $f(\cdot)$ and $g(\cdot)$ in the following text. The content in the latent feature of the source image is changing as the diffusion process continues. Therefore, the location of the correspondent pixel in latent space may change over diffusion steps. A direct solution is to decode the latent feature ${\boldsymbol{z}}$ into an image at each step and extract the mask dynamically according to the object location in the generated image. However, we find that a changing mask usually confuses the model and leads to a less optimal performance. Therefore, we use the same high-quality mask through the diffusion process. We find that the mask could either be extracted from the source image directly using an off-the-shelf model or from the generation process as in [37, 31]. Please check the Appendix for more details.
@@ 4.2.2 Style Adaptation
An ideal object swapping should keep the style unchanged. The object information in the generated variables is injected via the new concept token. Some style attributes could be already bound with the token. Therefore, solely generating the foreground information via the text prompt might lead to style inconsistency. Recently, [19, 30] found that adding such normalization layers can help improve the conditional image generation quality because such activation functions modulation. Unlike them, we employ the AdaIN (Adaptive Instance Normalization) to modulate the swapping features with spatial constraints. We follow Eq. 6 and Eq. 7 to denormalize the $V_{concept}$ with the mean and variance from $V_{src}$ in each time step for $V_{target}$ during the image generation process. As a result, we find that by modulating the concept feature, the generated content can adaptively follow the original style in the source image.
@@ 4.2.3 Scale Adaptation
Guidance from an external classifier in the inference process of Diffusion models could influence the diffusion noise to control the generated image. [10] shows that the guidance can also be used on the attention map to control the generation. Similarly, we adapt the mask guidance Eq. 9 to better align the shape between the source object and the target object.
@@
where $s$ is the classifier-free guidance strength and $v$ is an additional guidance weight for $g$. As with classifier guidance, we scale by $\sigma_{t}$ to convert the score function to a prediction of $\mathbf{\epsilon}_{t}$. $Shape(\mathbf{M_{src}}{})\left(k\right)$ denotes the object shape as identified in the cross-attention layer. Here the energy function $g$ is set as $\|\mathbf{M_{src}}{}-Shape(\mathbf{M_{src}}{})\left(k\right)\|_{1}$ to calculate the shape difference between the original object mask and the extracted shape of object token $k$ in the attention layer, which indicates the deviation between the ideal shape and shape during the diffusion process.
@@ 4.2.4 Content Adaptation
The dilation of the mask $\mathbf{M_{src}}{}$ using the structuring element K, where denotes the dilation operation and G is the Gaussian kernel. The asterisk * denotes the convolution operation. $S^{\prime}$ is the final soft mask.
@@ 5.1 Implementation Details
Here we introduce implementation details. In our paper, we used Stable Diffusion 2.1 as the pre-trained text-to-image diffusion model. DreamBooth [34] is used to convert the concept into textual space. We used null-text inversion [28] based on DDIM inversion [38] to boost the accuracy. There is no additional operation for single-object, partial object, cross domain swapping. Multi-object swapping is achieved by conducting swapping operation on the previous swapped image. Please check Appendix for evaluation dataset details.
Baselines setting. Photoswap[13], P2P[15], PnP[40], MasaCtrl[3] are attention variable based image editing methods, which are also compatible with our proposed Masked Latent Blending in Sec. 4.1 and Location Adaptation in Sec. 4.2. Therefore we boost their performance with those additional components. Otherwise, their performance would be much worse. Please see appendix for comparisons with their original methods, and the implementation details. Also, please check appendix for its performance on object insertion.
@@ 5.2 Single-object Swapping
We consider human evaluation to be the main quantitative performance measurement. A successful swap should keep the non-object area unchanged, change the object identity to target, and keep the gesture the same as the source object. As in Tab. 1, our model consistently outperforms baselines across all metrics. Fig. 4 shows the qualitative comparison for both human and non-human images. Thanks to the addition of our targeting variable swapping and location adaptation, all attention-manipulation based baselines also achieved perfect background preservation and some level of localized swapping result. SwapAnything yield a much better appearance adaptation result.
5.3 Multi-object Swapping
Figure 5: Multi-object swapping results of SwapAnything. Our method could easily swap multiple objects via swapping one object at a time. Note that the red circle means the target object to be replaced. The same color means a pair of concept and target for object swapping.
5.4 Partial Object Swapping
5.5 Cross-domain Swapping
Fig. 7 demonstrates that SwapAnything can adeptly handle a range of stylized source images, successfully adapting concept objects to match the desired style within the source image while seamlessly transferring identity into the generated images. For instance, when the source image is a photo of a certain style, yet SwapAnything skillfully generates the same painting style featuring personalities like “Charles Darwin” and “J. Robert Oppenheimer”. Notably, all the concept images are regular, unstyled photos, underscoring the model’s ability to blend different styles and identities effectively.
Figure 7: Results on cross-domain object swapping. With a variety of source images, including graphic, black-white photos, and oil paintings, the framework seamlessly integrates concept objects taken from regular images into these diverse source images.
5.6 Text-based Swapping
As shown in Fig. 8, besides personalized swapping, our method can also do text-based swapping, swapping an object in the source image with another described in text. This can be achieved by simply replacing the personalized concept token * with a text prompt, e.g., “A photo of $new\_obj$”.
5.7 Ablation Study
Fig. 9 show the effect of the components in SwapAnything. From the left part, we see that without latent feature swap, even with a mask and attention variable swap, the context pixel is still changed. Both latent feature and attention variable has effect of information preservation when compared with the result of no swap. The effect of adaptation and mask is presented on the right part. With style adaptation, the visual texture is closer to the source image. Without scale adaptation, the shape is not well aligned and artifacts appear in the neck part. When without any adaptation, the generated image is much less connected the source image regarding the swapping area. When without mask, both background and targeting area are changed, which leads to a different image.
Figure 9: Ablation study. The left part shows the effect of swapping, and the right part shows the effect of adaptation and mask.
5.8 Ethics Discussion
While SwapAnything achieves impressive performance across various visual tasks, it also raises potential ethical concerns. Like other advanced image editing technologies, SwapAnything could be misused to create deceptive content. To mitigate such risks, implementing technologies like digital watermarking can help track modifications and authenticate the sources of images, assisting users in distinguishing between genuine and altered content. Moreover, advocating for clear guidelines and regulations that govern the ethical use of image manipulation technologies is essential. These measures ensure responsible development and use, fostering a balance between innovation and ethical considerations. Here, we emphasize the importance of applications involving human face swapping. We strongly discourage using our methods on human faces without consent, and we are committed to preventing misinformation and harm to any individual or community. We enforce strict consent protocols in related tasks, requiring explicit permission from individuals before their images are used, especially in sensitive applications like face-swapping.
6 Conclusion
SwapAnything represents a notable breakthrough in the realm of object swapping. Swapping latent features and attention variables in the diffusion model ensures the retention of crucial information within the generated image. Through a targeted manipulation, we achieved a perfect background preservation. Additionally, we have introduced a sophisticated appearance adaptation process designed to seamlessly integrate the concept into the context of the source image. Consequently, SwapAnything is equipped to handle a diverse array of object swapping challenges. In the future, we plan to extend our framework to 3D/video personalized object swapping tasks.
Appendix 0.A Object Insertion
Appendix 0.B Variable Swapping Details
We use Stable Diffusion 2.1 as the pre-trained text-to-image diffusion model. DreamBooth [34] is used to convert the concept into textual space. The learning rate for this process is set at 1e-6, and we use the Adawm optimizer for 800 steps. The U-net and the text encoder are fine-tuned during this process, typically taking about 2 minutes on a machine equipped with 8 A100 GPUs. The target prompt is essentially the source prompt with a swap in object tokens to introduce a new concept.
For object mask, we first detect the object with Grounding DINO [25] and then extract the mask using Segment Anything [20]. For the targeting variable swapping, we do 30 for latent image feature $z$, 20 steps for cross-attention map, 25 for the self-attention map, and 10 for the self-attention output, we conduct swapping in all U-Net layer.
Appendix 0.C Adaptation Details
Scale Adaptation. We adapt the scale of the object in latent space to the shape of the mask. The object shape is indicated in the cross-attention map at each diffusion step [13, 15]. $Shape(\mathbf{M_{src}}{})(k)$ means the attention map for object text token $k$, which is obtained through binary-like transformation to the attention map. We apply a threshold of 0.4 after using sigmoid to normalize the attention value between 0 and 1.
Appendix 0.D Dataset
We conducted experiments on both human and non-human objects. For human swapping, we collect 50 faces as concepts. We also collected 500 images containing 1 or more people as the source images. For non-human object, we include DreamEdit [22] dataset and more concepts and its corresponding source images. In total, we aggregated 1,000 images.
Appendix 0.E Human Evaluation Interface
Figure 11: The illustration of the user study interface.
Appendix 0.F More Qualitative Results
Figure 12: Comparison on single-object swapping with baselines in their original components. Please zoom in for a clear visual result.
Here we first show the comparison with baselines in their original setting on single-object swapping. On other more challenging tasks, we also show the results of Photoswap since it is the state-of-the-art method of subject swapping. For the implementation details, we use external mask to help the inpainting process in DreamEdit. SwapAnything is also compared with BlipDiffusion [21]. Photoswap, P2P, PnP, and MasaCtrl, DreamEdit were equipped with the same DreamBooth model to grasp the new concept. Note that this would also indirectly include comparison with CustomEdit [6], since it also achieved personalized object swapping via equiping P2P with concept learning. CopyPaste involves directly transplanting the concept object in the concept image into the source object’s position.
Multi-object Swapping. Multi-object swapping is a big step after single-object swapping. First, previous methods usually have a background modification such that continuous editing would accumulate unwanted distortion, which leads to a totally different image and fails the task of swapping. The second issue is that previous methods are usually designed for main subject swapping and do not pay attention to other objects. In this case, the objects in the following swapping steps could disappear in the previous swapping process.
Appendix 0.G More Quantitative Results